Crane, Inc., has four-year bonds outstanding that pay a coupon rate of 6.50 percent and make coupon payments semiannually. If these bonds are currently selling at $917.89.

What is the yield to maturity that an investor can expect to earn on these bonds? (Round answer to 1 decimal place, e.g. 15.2%.)

Yield to maturity %


What is the effective annual yield? (Round answer to 1 decimal place, e.g. 15.2%.)

Effective annual yield
Answer #1

1)

Number of periods = 4 * 2 = 8

Coupon = (0.065 * 1000) / 2 =32.5

Yield to maturity = 9.0%

Keys to use in a financial calculator:

2nd I/Y 2

FV 1000

PMT 32.5

N 8

PV -917.89

CPT I/Y

2)

Effective annual yield = (1 + YTM/n)n - 1

Effective annual yield = (1 + 0.09/2)2 - 1

Effective annual yield = (1.045)2 - 1

Effective annual yield = 1.092 - 1

Effective annual yield = 0.092 or 9.2%